DEHYROGENATION OF ETHYLENE: SPECTROSCOPY AND STRUCTURES OF La(C_{2}H_{2}) AND La(C_{4}H_{6}) COMPLEXES

\underline{SUDESH KUMARI}, MOURAD ROUDJANE, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Hydrogen elimination is observed in the reaction of laser-ablated La atoms and ethylene (C_{2}H_{4}) in a pulsed molecular beam source. Dehydrogenated products, La(C_{2}H_{2}) and La(C_{4}H_{6}), are identified by time-of-flight mass spectrometry and studied by pulsed-field-ionization zero-electron kinetic energy spectroscopy and density functional theory calculations. La(C_{2}H_{2}) is determined as a triangle (C_{2v}) in the ^{2}A_{2} ground electronic state, where La binds with C_{2}H_{2} in a two-fold mode (η^{2}). La(C_{4}H_{6}) is identified as a diligand species with La being sandwiched between C_{2}H_{2} and C_{2}H_{4}, each in a two-fold binding mode, and the complex is in the ^{2}A_{1} ground electronic state. The adiabatic ionization energies of La(η^{2}-C_{2}H_{2}) and La(η^{2}-C_{2}H_{2})(η^{2}-C_{2}H_{4}) are measured to be 41174(5) and 39405(5) cm^{−1}, respectively. La^{+}-C_{2}H_{2} and La^{+}-C_{4}H_{6} stretching and C-H bending frequencies of the corresponding ions are also determined, and the vibrational assignments are confirmed with deuterated ethylene measurements.